SOLVE x²y - show Immal value problem Cauchy euler 3xy + 4yooy (A) 7,5 yC1)=3
Answer #1


or 22 sol. z?y - 3xy +44= 0 - (() let x=e? so that z = enca) Also xy. - D(D-1), xy= DY where D=d dz So that (DCD-1)- 30+4